OpenStudy (anonymous):

find the domain g(x) = 4/2-5x

OpenStudy (zarkon):

if the equation is \[\frac{4}{2-5x}\] then the domain is all reals except when the denominator is zero 2-5x=0 2=5x x=2/5 so all reals except 2/5